# Beyond The Rainbow: High Performance Deep Reinforcement Learning On A Desktop PC

### v1

@@ Appendix A Full Results Tables
Table A1: Maximum scores obtained during training (averaged over 100 episodes and all performed using random seeds) after 200M Frames on the Atari-60 benchmark. We compared against other non-recurrent algorithms where FE-Rainbow DQN refers to Fast and Efficient Rainbow DQN (Schmidt & Schmied, 2021), and M-IQN refers to Munchausen-IQN (Vieillard et al., 2020).
Game Random Human Rainbow DQN Rainbow DQN M-IQN FE-Rainbow DQN [^8] FE-Rainbow DQN uses Life Information, but only 10M frames. BTR
(Dopamine) (Full)
AirRaid 400 3000 14056 19111 53544
Alien 228 7128 3587 9492 4492 12508 19149
Amidar 6 1720 2630 5131 1875 2071 17808
Assault 222 742 3511 14198 7504 10710 19385
Asterix 210 8503 18367 428200 49865 346758 593650
Asteroids 719 47389 1489 2713 1685 12346 169272
Atlantis 12850 29028 838590 826660 918183 812826 899104
BankHeist 14 753 1148 1358 1292 1411 1598
BattleZone 2360 37188 40895 62010 52517 112652 168340
BeamRider 364 16926 6529 16850 12775 26399 138102
Berzerk 124 2630 842 2546 736 3388 6704
Bowling 23 161 49 30 32 41 48
Boxing 0 12 99 100 99 100 100
Breakout 2 30 120 418 320 538 677
Carnival 380 4000 5069 5588 6031
Centipede 2091 12017 6618 8167 4371 8368 76243
ChopperCommand 811 7388 12844 16654 4573 4208 980233
CrazyClimber 10780 35829 147743 168788 150783 140712 140723
DemonAttack 152 1971 17802 111185 68825 131657 135448
DoubleDunk -19 -16 22 0 22 -1 23
ElevatorAction 0 3000 79968 89237 82669
Enduro 0 860 2230 2126 2332 2266 2352
FishingDerby -92 -39 43 31 55 42 56
Freeway 0 30 34 34 34 34 34
Frostbite 65 4335 8572 9590 9538 5282 19332
Gopher 258 2412 10641 70355 27469 25607 99739
Gravitar 173 3351 1272 1419 1134 2107 5284
Hero 1027 30826 46764 55887 26037 15377 21560
IceHockey -11 1 2 1 12 7 38
Jamesbond 29 303 1106 1637 29828
JourneyEscape -18000 -1000 -959 -806 5166
Kangaroo 52 3035 13460 14638 10939 11499 13849
Krull 1598 2666 6229 8742 10703 10324 11124
KungFuMaster 258 22736 27900 52181 27119 27444 54330
MontezumaRevenge 0 4753 500 384 0 0 0
MsPacman 307 6952 4027 5380 6029 5982 11494
NameThisGame 2292 8049 9229 13136 12761 19819 28360
Phoenix 761 7243 8605 108529 5327 60954 350722
Pitfall -229 6464 -1 0 0 -2 0
Pong -21 15 20 21 19 21 21
Pooyan 500 1000 5640 13096 24280
PrivateEye 25 69571 21532 4234 100 254 100
Qbert 164 13455 18503 33818 14739 25712 39484
Riverraid 1338 17118 21091 16271 24585
RoadRunner 12 7845 55300 62041 61269 81832 590236
Robotank 2 12 66 61 73 71 83
Seaquest 68 42055 11362 15899 23885 63724 409991
Skiing -17098 -4337 -20518 -12958 -10336 -22077 -9132
Solaris 1236 12327 2438 3560 5765 2878 8198
SpaceInvaders 148 1669 4420 18789 13871 28099 53863
StarGunner 664 10250 57909 127029 65757 310404 574106
Tennis -24 -8 0 0 0 16 24
TimePilot 3568 5229 12283 12926 15155 31333 110981
Tutankham 11 168 245 241 207 167 315
UpNDown 533 11693 39065 216080 397876
Venture 0 1188 1579 6 1101 437 0
VideoPinball 0 17668 513484 533936 625118 269619 589066
WizardOfWor 564 4756 8201 17862 13644 15519 50828
YarsRevenge 3093 54577 45567 102557 111583 98908 177431
Zaxxon 32 9173 15089 22210 19080 18833 47096
IQM 0.000 1.000 1.994 2.718 2.609 2.687 7.419
Median 0.000 1.000 1.533 2.270 1.774 1.906 4.695
Mean 0.000 1.000 4.319 7.828 6.319 7.670 19.697
Best - - 1 7 5 3 50
>Human - - 45 37 41 38 52
@@
Defender 2875 18689 55105 169930 461380
Table A2: Maximum scores obtained during training (averaged over 100 episodes and all performed random seeds) after 200M Frames on the Atari-5 Environment, compared against other non-recurrent non-distributed algorithms. FE-Rainbow refers to Fast and Efficient Rainbow DQN (Schmidt & Schmied, 2021), and M-IQN refers to Munchausen-IQN (Vieillard et al., 2020).
@@
NameThisGame 2292 8049 9229 13136 12761 19819 28710
Phoenix 761 7243 8605 108529 5327 60955 367284
QBert 164 13455 18503 33818 14739 25712 45034
IQM 0.000 1.000 1.265 3.583 1.452 4.070 7.627
Median 0.000 1.000 1.21 2.532 1.44 3.167 4.589
Mean 0.000 1.000 3.714 5.817 3.745 4.684 16.561
Appendix B Full Results Graph
Figure B1: Performance of BTR on Atari-5 with individual components removed using Human-Normalized Interquartile Mean. Averaged over 3 seeds.
Figure B2: Performance of BTR on each individual game in the Atari-60 benchmark.
Figure B3: Performance of BTR on each individual game in the Procgen benchmark. Shaded areas show one standard deviation of the performed evaluations. The red dotted line shows the performance of Rainbow DQN + Impala with 4x scaled Impala blocks (Cobbe et al., 2020), after 200M frames.
Appendix C Hyperparameters
C.1 Environment Details
Table C3: Environment Details for Atari Experiments.
@@ C.1 Environment Details
Table C4: Environment Details for Procgen Experiments.
@@
Grey-Scaling True
@@
C.2 Algorithm Hyperparameters
Table C5: Table showing the hyperparameters used in the BTR algorithm.
@@ C.2 Algorithm Hyperparameters
$\epsilon$-greedy decay 2M Frames
@@
Adam Epsilon Parameter 1.95e-5 (equal to $\frac{0.05}{batchsize}$)
@@
C.3 Clarity of the terms Frames, Steps and Transitions
Throughout the Arcade Learning Environment’s history (ALE) (Bellemare et al., 2013; Machado et al., 2018), there have been many ambiguities around the terms: frames, steps and transitions, which are sometimes used interchangeably. Frames refer to the number of individual frames the agent plays, including those within repeated actions (also called frame skipping). This is notably different from the number of steps the agent takes, which does not include these skipped frames. When using the standard Atari wrapper, training for 200M frames is equivalent to training for 50M steps. Lastly, transitions refer to the standard tuple $(s_{t},a_{t},r_{t},s_{t+1})$, where the timestep $t$ refers to a steps, not frames. We encourage researchers to make this clear when publishing work, including when mentioning values of different hyperparameters.
Appendix D Beyond The Rainbow Architecture & Loss Function
D.1 Architecture
Figure D4 shows the the neural network architecture of the BTR algorithm. The architecture is highly similar to the Impala architecture (Espeholt et al., 2018), with notable exceptions:
Figure D4: Architectural diagram of the BTR algorithm’s neural network. The model contains a total of 2.91 million parameters, 2.52 million of which are within linear layers.
@@ D.1 Architecture
Within each Impala CNN blocks, each residual layer (containing two Conv 3x3 + ReLu) has spectral normalization applied, as discussed in Section 3.1.
@@
In order to use IQN, it is required to draw Tau samples which are multiplied by the output of the CNN layers, as shown by the section ‘IQN Samples’ in figure D4.
@@
Lastly, the sizes of many of the layers given in Figure D4 are dependant upon the Impala width scale, of which we use the value 2. For example, the Impala CNN blocks have [16$\times$×width, 32$\times$×width, 32$\times$×width] channels respectively. The output size of the convolutional layers (including the maxpooling layer) is 6$\times$6$\times$32$\times$×width, as a 6x6 maxpooling layer is used. Lastly, the cos embedding layer after generating IQN samples requires the same size as the output of the convolutional layers, hence the size is selected accordingly. Another benefit of the 6x6 maxpooling layer is following the product of the convolutional layers and IQN samples, the number of parameters is fixed, regardless of the input size. Figure D5 shows the numbers of parameters the ablated versions of BTR have.
Figure D5: Total number of parameters in BTR with different components removed. Those not included in the graph (Munchausen and Spectral Normalisation) used the same number of parameters as BTR.
D.2 Loss Function
@@
$TD_{BTR}=r_{t}+\alpha[\tau\ln\pi(a_{t}|s_{t})]_{l_{0}}^{0}+\gamma\sum_{a\in A}\pi(a|s_{t+1})(z_{\sigma^{{}^{\prime}}}(s_{t+1},a)-\tau\ln\pi(a|s_{t+1}))-z_{\sigma}(s_{t},a_{t})$ (D1)
@@
with $\pi(\cdot|s)=sm(\frac{\tilde{q}(s,\cdot)}{\tau})$ (that is, the policy is softmax with q˜, the quantity with respect to which the original policy of IQN is greedy). It is also worth noting here that due to the character conflict of both Munchausen and IQN using $\tau$ (Munchausen as a temperature parameter, and IQN for drawing samples), we replace IQN’s $\tau$ with $\sigma$. $l_{0},\tau$ and $\alpha$ are hyperparameters set by Munchausen. We use the same values in BTR, also shown in our hyperparameter table in Appendix C.2.
Appendix E BTR with Fewer Training Frames
Some of BTR’s improvements provide a relatively small improvement after 200M frames, however we want to point out their importance using fewer samples. Table E6 shows that many improvements provide large benefits earlier in training.
Table E6: A comparison of BTR’s ablations when using less than 200M frames on the Atari-5 benchmark. Percentages are given relative to BTR’s score.
Algorithm 40M Frames 80M Frames 120M Frames
BTR w/o Maxpooling 5.399 (+90%) 6.895 (+4%) 7.443 (+2%)
BTR 2.837 ($\pm$ 0%) 6.613 ($\pm$0%) 7.297 ($\pm$0%)
BTR w/o IQN 4.030 (+42%) 3.888 (-41%) 6.644 (-9%)
BTR w/o Spectral Normalisation 2.145 (-24%) 5.752 (-13%) 6.351 (-13%)
BTR w/o Munchausen 2.114 (-25%) 3.341 (-49%) 4.755 (-35%)
BTR w/o Impala 0.535 (-81%) 1.183 (-82%) 1.372 (-81%)
Appendix F Analysis of BTR’s Extensions
In recent years, several measures have been devised to understand the impact of extensions on a model’s behaviour, one of the most popular of which is dormant neurons (Kumar et al., 2020; Sokar et al., 2023). Figure F6 plots their prevalence for each of the ablations. We observe three key changes from the ablations; within the CNN layers, the Nature CNN (as used in Rainbow DQN) has a higher number of dormant/low activation neurons, indicating that the Impala network is significantly better for reducing dormant neurons. As for the fully connected layers, dormant neurons were very high across the board, with IQN lowering the number of dormant neurons.
(a) Convolutional Layers.
(b) Fully Connected Layers.
Figure F6: Histogram showing the percentage of average neuron activations for both convolutional and fully connected layers on Atari BattleZone, NameThisGame and Qbert, with subplots for each performed ablation. The first bar in each plot before the red dotted line represent neurons which Sokar et al. (2023) defined as dormant.
Appendix G BTR with and without Epsilon Greedy

### v2

@@ Appendix A Full Results Tables
Table A1: Maximum scores obtained during training (averaged over 100 episodes and all performed using random seeds) after 200M Frames on the Atari-60 benchmark. Fast & Efficient Rainbow DQN and Munchausen-IQN refer to Schmidt & Schmied (2021) and (Vieillard et al., 2020) respectively. FE-Rainbow uses Life Information (See Appendix I), only 10M frames, and has missing games, so metrics are based on existing games.
Game Random Human DQN (Nature) Rainbow M-IQN FE-Rainbow BTR
AirRaid 400 1000 7523 12472 19111 51719
Alien 227 7127 2354 3610 4249 12508 18999
Amidar 5 1719 1268 2390 1653 2071 14027
Assault 222 742 1526 3490 6014 10709 19064
Asterix 210 8503 2803 16547 42615 346758 608829
Asteroids 719 47388 846 1494 1666 12345 153589
Atlantis 12850 29028 843372 791393 866810 812825 891773
BankHeist 14 753 560 1070 1305 1411 1482
BattleZone 2360 37187 18425 40316 50501 112652 168340
BeamRider 363 16926 5203 6084 12322 26398 110415
Berzerk 123 2630 467 832 719 3388 11417
Bowling 23 160 30 43 23 40 63
Boxing 0 12 79 98 99 99 100
Breakout 1 30 92 109 241 537 682
Carnival 380 4000 5111 4523 5588 6284
Centipede 2090 12017 2378 6595 4425 8368 64242
ChopperCommand 811 7387 2722 13029 551 4208 956870
CrazyClimber 10780 35829 103549 146262 146419 140712 140927
DemonAttack 152 1971 5437 17411 63143 131657 135626
DoubleDunk -18 -16 -5 22 21 -1 23
ElevatorAction 0 3000 408 79372 89237 76941
Enduro 0 860 642 2165 2247 2266 2358
FishingDerby -91 -38 -1 42 54 42 58
Freeway 0 29 26 33 33 34 34
Frostbite 65 4334 482 8309 9419 5282 15158
Gopher 257 2412 5440 9987 23310 25606 97879
Gravitar 173 3351 209 1249 1105 2107 4253
Hero 1027 30826 15766 46290 25555 15377 25371
IceHockey -11 0 -6 0 11 6 44
Jamesbond 29 302 671 995 1526 59991
JourneyEscape -18000 -1000 -3300 -1096 -806 2841
Kangaroo 52 3035 10744 13005 10704 11498 14300
Krull 1598 2665 6029 4368 10309 10324 11268
KungFuMaster 258 22736 22397 27066 25588 27444 53302
MontezumaRevenge 0 4753 0 500 0 0 0
MsPacman 307 6951 3431 3989 5630 5981 13200
NameThisGame 2292 8049 7549 8900 12440 19819 27917
Phoenix 761 7242 4993 8800 5315 60954 427481
Pitfall -229 6463 -45 -27 -32 -1 0
Pong -20 14 16 20 19 21 21
Pooyan 500 1000 3452 4344 13096 22003
PrivateEye 24 69571 1113 21353 100 253 100
Qbert 163 13455 9801 18332 13159 25712 42927
Riverraid 1338 17118 9725 20675 16143 25192
RoadRunner 11 7845 38430 55104 60370 81831 579800
Robotank 2 11 59 67 71 70 83
Seaquest 68 42054 2416 9590 23885 63724 428263
Skiing -17098 -4336 -16281 -29268 -10404 -22076 -7938
Solaris 1236 12326 1478 1686 1835 2877 6301
SpaceInvaders 148 1668 1797 4455 10810 28098 54262
StarGunner 664 10250 48498 57255 64875 310403 577547
Tennis -23 -8 -3 0 0 15 24
TimePilot 3568 5229 3704 11959 14600 31333 113801
Tutankham 11 167 103 244 205 167 297
UpNDown 533 11693 8797 37936 197043 391439
Venture 0 1187 13 1537 978 437 1
VideoPinball 0 17667 38720 460245 508012 269619 573774
WizardOfWor 563 4756 1473 7952 11352 15518 47314
YarsRevenge 3092 54576 23963 46456 106929 98908 209499
Zaxxon 32 9173 4471 14983 14286 18832 52619
IQM ($\uparrow$) 0.000 1.000 0.771 1.852 2.181 $\approx 2.769$ 7.361
Median ($\uparrow$) 0.000 1.000 0.731 1.506 1.559 $\approx 1.906$ 4.690
Mean ($\uparrow$) 0.000 1.000 2.261 4.152 5.260 $\approx 7.700$ 21.574
Optimality Gap ($\downarrow$) 0.000 1.000 0.407 0.200 0.224 $\approx 0.180$ 0.098
Best - - 0 3 3 2 54
>Human - - 22 43 34 38 52
@@
Defender 2875 18689 169929 461380
Table A2: Maximum scores obtained during training (averaged over 100 episodes and all performed 3 random seeds) after 200M Frames on the Atari-5 Environment, compared against other non-recurrent non-distributed algorithms. FE-Rainbow refers to Fast and Efficient Rainbow DQN (Schmidt & Schmied, 2021), and M-IQN refers to Munchausen-IQN (Vieillard et al., 2020). Metrics do not use the recommended regression procedure, as explained in Appendix K.
@@
NameThisGame 2292 8049 9229 13136 12761 19819 27917
Phoenix 761 7243 8605 108529 5327 60955 427481
QBert 164 13455 18503 33818 14739 25712 42927
IQM 0.000 1.000 1.265 3.583 1.452 4.070 7.739
Median 0.000 1.000 1.21 2.532 1.44 3.167 4.766
Mean 0.000 1.000 3.714 5.817 3.745 4.684 18.453
Table A3: Comparison of performance and walltime against PQN (Gallici et al., 2024). PQN only reports results at 400M frames and includes life information, which greatly affects performance (see Appendix I). To provide a fairer comparison, we also report our results using life information but only use 200M frames. Below are Atari-5 IQM and per-game Scores, with BTR averaged over 3 seeds. Human-Normalized scores are reported for individual games, with the raw score in brackets.
Game BTR (with life info, 200M frames) PQN (with life info, 400M frames)
Inter-Quartile Mean 12.18 3.86
BattleZone 12.73 (445,827) 1.51 (54,791)
DoubleDunk 14 (23.0) 6.03 (-0.92)
NameThisGame 4.51 (28,834) 3.18 (20,603)
Phoenix 90.85 (589,662) 38.79 (252,173)
QBert 9.79 (130,348) 2.37 (31,716)
Walltime (A100) 22 Hours 2 Hours
Backend (PyTorch (non-compiled) + gymnasium async) (JAX + envpool)
Appendix B Full Results Graphs
Figure B1: Performance of BTR (4 Seeds) on each individual game in all 60 Atari games. Shaded areas show 95% confidence intervals over different seeds.
Figure B2: Final performance of BTR on Atari-60 (as used in RLiable (Agarwal et al., 2021)), against other popular algorithms. The plot displays performance profiles, with 95% confidence intervals and 4 seeds for BTR.
Figure B3: Performance of BTR with different components removed on individual games, and interquartile mean in the top left. Results are averaged over 4 seeds, with shaded areas showing 95% bootstrapped confidence intervals.
Figure B4: Figure shows BTR against PPO on Atari 57 (PPO’s Atari 60 scores have not been reported). PPO uses the results provided by DreamerV3 (Hafner et al., 2023), which additionally uses vectorization and the impala algorithm. Shaded areas show 95% confidence intervals with BTR using 4 seeds.
Figure B5: Performance of BTR on each individual game in the Procgen benchmark. Shaded areas show 95% confidence intervals. The red dotted line shows the performance of Rainbow DQN + Impala with 4x scaled Impala blocks (Cobbe et al., 2020), after 200M frames.
Appendix C Additional Ablations
Figure C6: Figures show BTR’s human-normalized scores without different components, with shaded areas showing 95% bootstrapped confidence intervals averaged over 4 seeds. Left: Predicted Atari-60 median score using the regression procedure defined in Aitchison et al. (2023). However, we find the predicted median does not match the true median (see Appendix K). Right: Interquartile mean across the 5 games.
Appendix D Hyperparameters
D.1 Environment Details
Table D4: Environment Details for Atari Experiments.
@@ D.1 Environment Details
Table D5: Environment Details for Procgen Experiments.
@@
Grey-Scaling False
@@
D.2 Algorithm Hyperparameters
Table D6: Table showing the hyperparameters used in the BTR algorithm.
@@ D.2 Algorithm Hyperparameters
$\epsilon$-greedy Decay 8M Frames
@@
Adam Epsilon Parameter 1.95e-5 (equal to $\frac{0.005}{batchsize}$)
@@
D.3 Clarity of the terms Frames, Steps and Transitions
Throughout the Arcade Learning Environment’s history (ALE) (Bellemare et al., 2013; Machado et al., 2018), there have been many ambiguities around the terms: frames, steps and transitions, which are sometimes used interchangeably. Frames refer to the number of individual frames the agent plays, including those within repeated actions (also called frame skipping). This is notably different from the number of steps the agent takes, which does not include these skipped frames. When using the standard Atari wrapper, training for 200M frames is equivalent to training for 50M steps. Lastly, transitions refer to the standard tuple $(s_{t},a_{t},r_{t},s_{t+1})$, where the timestep $t$ refers to a steps, not frames. We encourage researchers to make this clear when publishing work, including when mentioning the values of different hyperparameters.
Appendix E Beyond The Rainbow Architecture & Loss Function
E.1 Architecture
Figure E7 shows the neural network architecture of the BTR algorithm. The architecture is highly similar to the Impala architecture (Espeholt et al., 2018), with notable exceptions:
Figure E7: Architectural diagram of the BTR algorithm’s neural network. The model contains a total of 2.91 million parameters, 2.52 million of which are within linear layers.
@@ E.1 Architecture
Within each Impala CNN block, each residual layer (containing two Conv 3x3 + ReLu) has spectral normalization applied, as discussed in Section 3.1.
@@
In order to use IQN, it is required to draw Tau samples, which are multiplied by the output of the CNN layers, as shown by the section ‘IQN Samples’ in figure E7.
@@
Lastly, the sizes of many of the layers given in Figure E7 are dependent upon the Impala width scale, of which we use the value 2. For example, the Impala CNN blocks have [16$\times$×width, 32$\times$×width, 32$\times$×width] channels respectively. The output size of the convolutional layers (including the maxpooling layer) is 6$\times$6$\times$32$\times$×width, as a 6x6 maxpooling layer is used. Lastly, the cos embedding layer, after generating IQN samples, requires the same size as the output of the convolutional layers. Hence, the size is selected accordingly. Another benefit of the 6x6 maxpooling layer is following the product of the convolutional layers and IQN samples, the number of parameters is fixed, regardless of the input size. Figure E8 shows the number of parameters the ablated versions of BTR have.
Figure E8: Total number of parameters in BTR with different components removed. Those not included in the graph (Munchausen and Spectral Normalization) used the same number of parameters as BTR.
E.2 Loss Function
@@
$TD_{BTR}=r_{t}+\alpha[\tau\ln\pi(a_{t}|s_{t})]_{l_{0}}^{0}+\gamma\sum_{a\in A}\pi(a|s_{t+1})(z_{\sigma^{{}^{\prime}}}(s_{t+1},a)-\tau\ln\pi(a|s_{t+1}))-z_{\sigma}(s_{t},a_{t})$ (E1)
@@
with $\pi(\cdot|s)=sm(\frac{\tilde{q}(s,\cdot)}{\tau})$ (that is, the policy is softmax with q˜, the quantity with respect to which the original policy of IQN is greedy). It is also worth noting here that due to the character conflict of both Munchausen and IQN using $\tau$ (Munchausen as a temperature parameter, and IQN for drawing samples), we replace IQN’s $\tau$ with $\sigma$. $l_{0},\tau$ and $\alpha$ are hyperparameters set by Munchausen. We use the same values in BTR, also shown in our hyperparameter table in Appendix D.2.
E.3 Analysis Confidence Intervals
Due to space constraints, we was unable to include confidence intervals for Table 2 in the main paper. A repeat of the main paper Table can be found in Table E7, with the associated confidence intervals in Table E8.
Table E7: Repeat of the main paper’s Table 2 for reference against the table of 95% confidence intervals below. Comparison of policy churn, action gaps, actions swaps and evaluation performance with different quantities of $\epsilon$-actions and color jitter (both only applied for evaluation). All measurements use the final agent, trained on 200 million frames, for Atari Phoenix, averaged over 3 seeds. Action Gap is the average absolute Q-value difference between the highest two valued actions. % Actions Swap is the percentage of times the agent’s argmax action has changed from the last timestep. Policy churn is the percentage of states which the agent’s argmax action has changed on after a single gradient step. Color jitter applies a random 10% change to the brightness, saturation and hue of each frame. For associated error with these values, please see Appendix E.3.
Category BTR w/o Munchausen w/o IQN w/o SN w/o Impala w/o Maxpool
Action Gap 0.282 0.055 0.180 0.274 0.215 0.264
% Action Swaps 36.6% 47.7% 42.2% 40.3% 41.1% 39.3%
Policy Churn 3.8% 11.0% 0.5% 3.3% 4.5% 4.2%
Score ColorJitter 212k 85k 110k 187k 19k 187k
Score $\epsilon=0.03$ 94k 42k 62k 75k 10k 86k
Score $\epsilon=0.01$ 194k 70k 110k 132k 13k 171k
Score $\epsilon=0$ 330k 184k 187k 296k 21k 406k
Table E8: 95% confidence intervals for the main paper Table 2, calculated using 6 seeds. A repeat of the original table is shown above in Table E7.
Category BTR w/o Munchausen w/o IQN w/o SN w/o Impala w/o Maxpool
Action Gap [0.25, 0.31] [0.05, 0.06] [0.17, 0.19] [0.25, 0.3] [0.1, 0.33] [0.23, 0.3]
% Action Swaps [33.6, 39.6] [45.8, 49.7] [41.0, 43.3] [38.5, 42.2] [31.4, 50.8] [38.2, 40.4]
Policy Churn [3.3, 4.2] [10.2, 11.7] [0.5, 0.5] [2.9, 3.6] [3.9, 5.0] [3.7, 4.7]
Score ColorJitter [204k, 218k] [77k, 92k] [90k, 128k] [173k, 201k] [0k, 40k] [158k, 215k]
Score $\epsilon=0.03$ [88k, 99k] [38k, 46k] [52k, 72k] [68k, 82k] [2k, 18k] [74k, 97k]
Score $\epsilon=0.01$ [177k, 211k] [62k, 77k] [97k, 122k] [139k, 176k] [0k, 52k] [143k, 199k]
Score $\epsilon=0$ [282k, 377k] [116k, 251k] [163k, 209k] [244k, 348k] [0k, 43k] [332k, 479k]
Appendix F BTR with and without Epsilon Greedy